suppose 82% o all students aking a beginning p oaramming course fail to get her first program to run on firs subr his on. ConPlease solve and explain steps
Answer #1

a.)n=5 p=0.82

b.)

given P(X=x)=\binom{n}{x}p^x(1-p)^n^-^x

given P(X=4)=\binom{5}{4}0.82^4(1-0.82)^5^-^4

  =0.4069 =0.407

c.)

P(X\leqslant 4)=1-P(X=5))

  =1-\binom{5}{5}0.82^5(1-0.82)^5^-^5

=1-0.3707

=0.629

d.)

P(X\geqslant 4)=P(X=4)+P(X=5)

==\binom{5}{4}0.82^4(1-0.82)^5^-^4+\binom{5}{5}0.82^5(1-0.82)^5^-^5

  =0.7776=0.778

e.)

E(x)=np

=5*0.82

  =4.1

f)

(\sigma )x=\sqrt{n*p*(1-p)}

  (\sigma )x=\sqrt{5*0.82*0.18}

==0.859